# AttackSeqBench: Benchmarking Large Language Models’ Understanding of Sequential Patterns in Cyber Attacks

### v1

@@ C.1. Question Generation Prompts
[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic, technique and procedures given below.
Table 15. Few-shot prompt for question generation for AttackSeq-Procedure-No.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **Reference Question-Answer Pair** that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always ”Yes”. 2. A **Reference MITRE TTP** that is NOT supported in the CTI report. [End of Inputs]
Your task is to generate two questions based on the given reference question, such that attack sequence described in the question is modified and the correct answer to the two questions is ”No”. The definitions of the two questions are as follows: 1. Question 1 should negate the ”before” and/or ”after” clauses of the reference question, such that the attack sequence contradicts the original sequence of TTPs in the reference question. 2. Question 2 should replace the main procedures in the reference question with the provided reference MITRE TTP, such that the replaced main procedures are not found in the CTI report. [Steps]: Please follow these steps: 1. Analyze the Reference Question-Answer Pair: - Identify and outline the attack sequence in the order presented in the reference Question-Answer Pair. 2. Construct the Questions: - Design two questions that modify the attack sequence described in the reference question. - Ensure that the answer to the questions is ”No”. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the examples below for guidance. [Examples]: Example Questions: - Example 1: Reference Question: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely run commands on the command-line interface before achieving persistence by adding a Run key? Reference Answer: Yes Reference TTP: Tactic: Initial Access, Technique: T1651-Cloud Administration Command, Example Procedures: AADInternals can execute commands on Azure virtual machines using the VM agent. APT29 has used Azure Run Command and Azure Admin-on-Behalf-of (AOBO) to execute code on virtual machines. Pacu can run commands on EC2 instances using AWS Systems Manager Run Command. Question 1: After achieving persistence by adding a Run key, will the Ke3chang malware run commands on the command-line interface only after gaining initial access through compromised VPN accounts? Answer: No Question 2: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely execute commands on Azure virtual machines using the VM agent before achieving persistence by adding a Run key? Answer: No - Example 2: Reference Question: Will Axiom acquire dynamic DNS services for use in the targeting of intended victims before gaining initial access to the victim’s network using SQL injection? Reference Answer: Yes Reference TTP: Tactic: Resource Development, Technique: T1585.001-Social Media Accounts, Example Procedures: APT32 has set up Facebook pages in tandem with fake websites. Cleaver has created fake LinkedIn profiles that included profile photos, details, and connections. EXOTIC LILY has established social media profiles to mimic employees of targeted companies. Question 1: Will Axiom acquire dynamic DNS services for use in the targeting of intended victims only after gaining initial access to the victim’s network using SQL injection? Reference Answer: No Question 2: Will Axiom set up Facebook pages in tandem with fake websites before gaining initial access to the victim’s network using SQL injection? Answer: No - Example 3: Reference Question: Will Ke3chang perform frequent and scheduled data exfiltration from compromised networks after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Reference Answer: Yes Reference TTP: Tactic: Exfiltration, Technique: T1030-Data Transfer Size Limits, Example Procedures: AppleSeed has divided files if the size is 0x1000000 bytes or more. APT28 has split archived exfiltration files into chunks smaller than 1MB. APT41 transfers post-exploitation files dividing the payload into fixed-size chunks to evade detection. Question 1: Will Ke3chang perform frequent and scheduled data exfiltration from compromised networks only before establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: No Question 2: Will Ke3chang divide files if the size is 0x1000000 bytes or more after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: No [End of Examples]
3. Provide the Question-Answer Pairs: - Please follow the output format: ”Question 1: ¡insert question 1 here¿ Answer: ¡insert answer to question 1 here¿ Question 2: ¡insert question 2 here¿ Answer: ¡insert answer to question 2 here¿”.
[End of Steps] Following the steps above, please generate two questions based on the Reference Question-Answer Pair and Reference MITRE TTP given below. Please only provide the final output of the two questions.
C.2. Dataset Refinement Prompts
The prompt templates to filter based on the Answerability criteria is in Table 16, while the feedback and refinement prompts are in Table 17 and Table 18 respectively.
Table 16. Prompt template for verifying Answerability during Self-Refinement.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **TTP Description**: A reference description of the correct answer corresponding to the question. 3. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. [End of Inputs] Your task is to evaluate the answerability of the given question using the provided information in the CTI Outline and TTP description. We define answerability based on three factors below: 1. The correct answer must be supported by the CTI outline. 2. The correct answer must clearly stand out as the best answer choice to the question based on the CTI outline. 3. Suppose the masked_tactic paragraph is removed from the CTI outline, the correct answer must be deducible from the answer choices by using the information provided in remaining tactics of the CTI outline and TTP description. You may also refer to your external cybersecurity knowledge to determine if the correct answer is deducible. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Match the correct answer with the provided TTP description. - Determine step-by-step if the question is answerable based on the definition above. 3. Output evaluation result: - Output one of the following: - ”A”: Indicates that the question is answerable. - ”B”: Indicates that the question is not answerable. - ”C”: Indicates that you do not know/cannot determine if the question is answerable. - Please also include a short and concise explanation of your evaluation result. - Please follow the output format: ”Explanation: ¡insert explanation here¿ Evaluation Result: ¡insert letter here¿”. [End of Steps] Following the steps above, please evaluate the question using the CTI report and description below and only output the evaluation result.
Table 17. Prompt template for feedback on question quality based on evaluation criteria.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the Tactics, Techniques and Procedures (TTPs) in MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **TTP Description**: A reference description of the correct answer to the question. 3. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI outline, along with one correct answer and distractors among the answer choices. [End of Inputs]
[Evaluation Criteria]: Your task is to evaluate the QA pair and provide your feedback for each of the criteria defined below: 1. **Clarity**: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. **Logical**: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. **Relevance**: Does the question directly relate to the CTI outline? 4. **Consistency**: Does the question align with the provided TTP Description? 5. **Answer Consistency**: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI outline: - Read the CTI outline carefully. - Identify and outline the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Assess each criterion step by step, rating it on a scale of 1 to 5 (1 = poor, 5 = excellent). - Provide a short and concise feedback for each rating. 4. Output Feedback Scores: - Please follow the output format: Feedback Scores: - Clarity: ¡Your feedback¿ (¡Score¿/5) - Logical: ¡Your feedback¿ (¡Score¿/5) - Relevance: ¡Your feedback¿ (¡Score¿/5) - Consistency: ¡Your feedback¿ (¡Score¿/5) - Answer Consistency: ¡Your feedback¿ (¡Score¿/5) Total Score: ¡Total Score¿/25 [End of Steps] Following the steps above, please evaluate the Question with Answer Choices below using the provided CTI report and TTP Description. Please only output the Feedback Scores.
Table 18. Prompt template for question refinement.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. **Feedback Results**: A list of feedback scores and explanations for each desired criterion of the question defined below. [End of Inputs] [Evaluation Criteria]: Please refer to the definition of each feedback criterion: 1. **Clarity**: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. **Logical**: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. **Relevance**: Does the question directly relate to the CTI outline? 4. **Consistency**: Does the question align with the provided TTP Description? 5. **Answer Consistency**: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? Your task is to iteratively refine the quality of the given question based on the feedback provided in the Feedback Results. [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. 2. Analyze the Question with Answer Choices: - Read the question and the provided answer choices carefully. 3. Analyze the Feedback Results: - Based on the feedback given in each criterion, refine the question to improve the each aspect. - Please ensure that the correct answer to the refined question is the same as the original question. - Please also ensure that the question avoids hinting at the correct answer. 4. Output the Refined Question: - Please follow the output format: ”Refined Question: ¡Your refined question here¿”. [End of Steps] Following the steps above, please refine the question based on the Feedback Results and CTI Outline provided below. Please only output the refined question.
C.3. Automatic Evaluation Prompt
We utilize the definitions in the evaluation criteria in Table 7 to create prompts. We show an example prompt template for evaluating the Logical aspect of the question shown in Table 19.
Table 19. Prompt template for evaluating the Logical aspect.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. **Description to Correct Answer**: The description of the correct answer from the MITRE ATT&CK framework. [End of Inputs] Your task is to rate the question on one metric below. [Definition of Logical Aspect] Evaluation Criteria: Logical (1-5): Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding and/or subsequent tactics in the CTI outline such that it logically leads to the correct answer? The scale is defined as follows: 1 - Not Logical: The question does not align with the logical sequence of MITRE ATT&CK tactics in the CTI outline. It ignores or contradicts the natural order of tactics and TTPs. 2 - Weak Logical Alignment: The question shows minimal alignment with the MITRE ATT&CK sequence. It may reference unrelated tactics or disrupt the logical flow. 3 - Moderately Logical: The question has some logical alignment, but it may not reference preceding or subsequent tactics clearly. The sequence could be improved. 4 - Strong Logical Alignment: The question follows the expected sequence of MITRE ATT&CK tactics and references preceding or subsequent TTPs in a logical manner. 5 - Perfect Logical Alignment: The question perfectly aligns with the MITRE ATT&CK framework, referencing relevant TTPs in a way that naturally leads to the correct answer. [Steps:] Evaluation Steps: 1. Analyze the CTI report and Description to Correct Answer: - Read the report and the provided description carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Evaluate the Question: - Read the question and the provided answer choices carefully. - Using the CTI outline and provided description to the correct answer, Rate the question on a scale of 1-5 according to the evaluation criteria above. 3. Output evaluation score: - Please only output the numerical evaluation score based on the defined criteria. [End of Steps] Following the steps above, please evaluate the question and only output the numerical evaluation score.
C.4. Answering Prompts
The prompt templates for the three benchmarking tasks, Regular, Zero-Shot and RAG-empowered are shown in Tables 20, 21, 22 respectively.
Table 20. Prompt template for Regular setting.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **Question** about a sequence of TTPs with several answer choices. [End of Inputs]
Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the CTI report. Note that the CTI report contains key details required for your analysis, but it may not directly state the answer. Your evaluation of the answer choices is essential to arrive at the correct answer. [Steps:] Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the Question: - Read the question and its answer choices. - Identify the sequence of TTPs mentioned in the question. 3. Compare and Evaluate: - Match the extracted attack sequence from the CTI report with the details in the question. - Evaluate each answer choice to determine which one aligns best with the attack sequence and any critical contextual information. 4. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”. [End of Steps] Following the steps above, please answer the question below using the provided CTI report.
Table 21. Prompt template for Zero-Shot setting.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive a question about a sequence of TTPs with several answer choices. Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the question. [End of Inputs]
[Steps:] Please follow these steps: 1. Analyze the Question: - Read the question and its answer choices. - Identify the sequence of TTPs mentioned in the question. 2. Compare and Evaluate: - Evaluate each answer choice to determine which one aligns best with the attack sequence in the question. 3. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”. [End of Steps] Following the steps above, please answer the question below.
Table 22. Prompt template for RAG-empowered setting.
@@ C.4. Answering Prompts
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **Question** about a sequence of TTPs with several answer choices. 2. A list of **Related TTPs** that are relevant to the question. [End of Inputs]
Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the question.

### v2

@@ 3. Provide the Question-Answer Pair:
- Please follow the output format: “Question: <insert question here> Answer: <insert answer here>”. [End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic, technique and procedures given below.
@@
Box 4: Few-shot prompt for question generation for AttackSeq-Procedure-No. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]:
You will receive two parts:
1. A Reference Question-Answer Pair that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always ”Yes”. 2. A Reference MITRE TTP that is NOT supported in the CTI report. [End of Inputs]
Your task is to generate two questions based on the given reference question, such that attack sequence described in the question is modified and the correct answer to the two questions is ”No”. The definitions of the two questions are as follows: 1. Question 1 should negate the ”before” and/or ”after” clauses of the reference question, such that the attack sequence contradicts the original sequence of TTPs in the reference question. 2. Question 2 should replace the main procedures in the reference question with the provided reference MITRE TTP, such that the replaced main procedures are not found in the CTI report. [Steps]:
Please follow these steps:
1. Analyze the Reference Question-Answer Pair: - Identify and outline the attack sequence in the order presented in the reference Question-Answer Pair. 2. Construct the Questions: - Design two questions that modify the attack sequence described in the reference question. - Ensure that the answer to the questions is ”No”. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the examples below for guidance. [Examples]:
Example Questions:
- Example 1: Reference Question: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely run commands on the command-line interface before achieving persistence by adding a Run key? Reference Answer: Yes Reference TTP: Tactic: Initial Access, Technique: T1651-Cloud Administration Command, Example Procedures: AADInternals can execute commands on Azure virtual machines using the VM agent. APT29 has used Azure Run Command and Azure Admin-on-Behalf-of (AOBO) to execute code on virtual machines. Pacu can run commands on EC2 instances using AWS Systems Manager Run Command. Question 1: After achieving persistence by adding a Run key, will the Ke3chang malware run commands on the command-line interface only after gaining initial access through compromised VPN accounts? Answer: No Question 2: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely execute commands on Azure virtual machines using the VM agent before achieving persistence by adding a Run key? Answer: No - Example 2: Reference Question: Will Axiom acquire dynamic DNS services for use in the targeting of intended victims before gaining initial access to the victim’s network using SQL injection? Reference Answer: Yes Reference TTP: Tactic: Resource Development, Technique: T1585.001-Social Media Accounts, Example Procedures: APT32 has set up Facebook pages in tandem with fake websites. Cleaver has created fake LinkedIn profiles that included profile photos, details, and connections. EXOTIC LILY has established social media profiles to mimic employees of targeted companies. Question 1: Will Axiom acquire dynamic DNS services for use in the targeting of intended victims only after gaining initial access to the victim’s network using SQL injection? Reference Answer: No Question 2: Will Axiom set up Facebook pages in tandem with fake websites before gaining initial access to the victim’s network using SQL injection? Answer: No - Example 3: Reference Question: Will Ke3chang perform frequent and scheduled data exfiltration from compromised networks after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Reference Answer: Yes Reference TTP: Tactic: Exfiltration, Technique: T1030-Data Transfer Size Limits, Example Procedures: AppleSeed has divided files if the size is 0x1000000 bytes or more. APT28 has split archived exfiltration files into chunks smaller than 1MB. APT41 transfers post-exploitation files dividing the payload into fixed-size chunks to evade detection. Question 1: Will Ke3chang perform frequent and scheduled data exfiltration from compromised networks only before establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: No Question 2: Will Ke3chang divide files if the size is 0x1000000 bytes or more after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: No [End of Examples]
3. Provide the Question-Answer Pairs:
- Please follow the output format: “Question 1: <insert question 1 here> Answer: <insert answer to question 1 here>.” “Question 2: <insert question 2 here> Answer: <insert answer to question 2 here>.” [End of Steps] Following the steps above, please generate two questions based on the Reference Question-Answer Pair and Reference MITRE TTP given below. Please only provide the final output of the two questions.
C.2 Dataset Refinement Prompt Templates
The prompt templates to filter based on the Answerability criteria is in Box C.2, while the feedback and refinement prompts are in Box C.2 and Box C.2 respectively.
Box 5: Prompt template for verifying Answerability during Self-Refinement. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. TTP Description: A reference description of the correct answer corresponding to the question. 3. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. [End of Inputs] Your task is to evaluate the answerability of the given question using the provided information in the CTI Outline and TTP description. We define answerability based on three factors below: 1. The correct answer must be supported by the CTI outline. 2. The correct answer must clearly stand out as the best answer choice to the question based on the CTI outline. 3. Suppose the masked_tactic paragraph is removed from the CTI outline, the correct answer must be deducible from the answer choices by using the information provided in remaining tactics of the CTI outline and TTP description. You may also refer to your external cybersecurity knowledge to determine if the correct answer is deducible. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Match the correct answer with the provided TTP description. - Determine step-by-step if the question is answerable based on the definition above. 3. Output evaluation result: - Output one of the following: - “A”: Indicates that the question is answerable. - “B”: Indicates that the question is not answerable. - “C”: Indicates that you do not know/cannot determine if the question is answerable. - Please also include a short and concise explanation of your evaluation result. - Please follow the output format: “Explanation: <insert explanation here> Evaluation Result: <insert letter here>.” [End of Steps]
Following the steps above, please evaluate the question using the CTI report and description below and only output the evaluation result.
Box 6: Prompt template for assessing question quality based on the evaluation criteria. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the Tactics, Techniques and Procedures (TTPs) in MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. TTP Description: A reference description of the correct answer to the question. 3. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI outline, along with one correct answer and distractors among the answer choices. [End of Inputs]
Your task is to evaluate the QA pair and provide your feedback for each of the criteria defined below: [Evaluation Criteria]:
Please refer to the definition of each feedback criterion:
1. Clarity: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. Logical: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. Relevance: Does the question directly relate to the CTI outline? 4. Consistency: Does the question align with the provided TTP Description? 5. Answer Consistency: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI outline: - Read the CTI outline carefully. - Identify and outline the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Assess each criterion step by step, rating it on a scale of 1 to 5 (1 = poor, 5 = excellent). - Provide a short and concise feedback for each rating. 4. Output Feedback Scores: - Please follow the output format: Feedback Scores: - Clarity: <Your feedback> (<Score>/5) - Logical: <Your feedback> (<Score>/5) - Relevance: <Your feedback> (<Score>/5) - Consistency: <Your feedback> (<Score>/5) - Answer Consistency: <Your feedback> (<Score>/5) Total Score: <Total Score>/25 [End of Steps] Following the steps above, please evaluate the Question with Answer Choices below using the provided CTI report and TTP Description. Please only output the Feedback Scores.
Box 7: Prompt template for question refinement. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. Feedback Results: A list of feedback scores and explanations for each desired criterion of the question defined below. [End of Inputs] Your task is to iteratively refine the quality of the given question based on the feedback provided in the Feedback Results. [Evaluation Criteria]: Please refer to the definition of each feedback criterion: 1. Clarity: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. Logical: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. Relevance: Does the question directly relate to the CTI outline? 4. Consistency: Does the question align with the provided TTP Description? 5. Answer Consistency: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. 2. Analyze the Question with Answer Choices: - Read the question and the provided answer choices carefully. 3. Analyze the Feedback Results: - Based on the feedback given in each criterion, refine the question to improve the each aspect. - Please ensure that the correct answer to the refined question is the same as the original question. - Please also ensure that the question avoids hinting at the correct answer. 4. Output the Refined Question: - Please follow the output format: “Refined Question: <Your refined question here>.” [End of Steps] Following the steps above, please refine the question based on the Feedback Results and CTI Outline provided below. Please only output the refined question.
C.3 Automatic Evaluation Prompt Templates
We utilize the definitions in the evaluation criteria in Table LABEL:tab.annotation_instructions to create prompts. We show an example prompt template for evaluating the Logical aspect of the question shown in Box C.3.
Box 8: Prompt template for evaluating the Logical aspect. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. Description to Correct Answer: The description of the correct answer from the MITRE ATT&CK framework. [End of Inputs] Your task is to rate the question on one metric below. [Definition]
Evaluation Criteria:
Logical (1-5): Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding and/or subsequent tactics in the CTI outline such that it logically leads to the correct answer? The scale is defined as follows: 1 - Not Logical: The question does not align with the logical sequence of MITRE ATT&CK tactics in the CTI outline. It ignores or contradicts the natural order of tactics and TTPs. 2 - Weak Logical Alignment: The question shows minimal alignment with the MITRE ATT&CK sequence. It may reference unrelated tactics or disrupt the logical flow. 3 - Moderately Logical: The question has some logical alignment, but it may not reference preceding or subsequent tactics clearly. The sequence could be improved. 4 - Strong Logical Alignment: The question follows the expected sequence of MITRE ATT&CK tactics and references preceding or subsequent TTPs in a logical manner. 5 - Perfect Logical Alignment: The question perfectly aligns with the MITRE ATT&CK framework, referencing relevant TTPs in a way that naturally leads to the correct answer. [End of the Definition]
[Steps:] Evaluation Steps: 1. Analyze the CTI report and Description to Correct Answer: - Read the report and the provided description carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Evaluate the Question: - Read the question and the provided answer choices carefully. - Using the CTI outline and provided description to the correct answer, Rate the question on a scale of 1-5 according to the evaluation criteria above. 3. Output evaluation score: - Please only output the numerical evaluation score based on the defined criteria. [End of Steps] Following the steps above, please evaluate the question and only output the numerical evaluation score.
C.4 Answering Prompt Templates
The prompt templates for the three benchmark settings (i.e., Context setting, Zero-Shot setting and RAG-empowered setting) are shown in Box C.4, Box C.4, Box C.4 respectively.
Box 9: Prompt template for the Context setting. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A Question about a sequence of TTPs with several answer choices. [End of Inputs]